### v1

# Can xLLMs Understand the Structure of Dialog? Exploring Multilingual Response Generation in Complex Scenarios

Abstract
Multilingual research has garnered increasing attention, especially in the domain of dialogue systems. The rapid advancements in large language models (LLMs) have fueled the demand for high-performing multilingual models. However, two major challenges persist: the scarcity of high-quality multilingual datasets and the limited complexity of existing datasets in capturing realistic dialogue scenarios. To address these gaps, we introduce XMP, a high-quality parallel Multilingual(X) dataset sourced from Multi-party Podcast dialogues [^1] https://github.com/ahxchzt/XMP_data. Each sample in the dataset features at least three participants discussing a wide range of topics, including society, culture, politics, and entertainment. Through extensive experiments, we uncover significant limitations in previously recognized multilingual capabilities of LLMs when applied to such complex dialogue scenarios. For instance, the widely accepted multilingual complementary ability of LLMs is notably impacted. By conducting further experiments, we explore the mechanisms of LLMs in multilingual environments from multiple perspectives, shedding new light on their performance in real-world, diverse conversational contexts.
Can xLLMs Understand the Structure of Dialog? Exploring Multilingual Response Generation in Complex Scenarios
Zhongtian Hu1, Yiwen Cui1, Ronghan Li2, Meng Zhao3, Lifang Wang1, 1School of Computer Science and Engineering, Northwestern Polytechnical University, 2School of Computer Science and Technology, Xidian University, 3School of Artificial Intelligence and Big Data, Henan University of Technology ahxchzt@mail.nwpu.edu.cn, wanglf@nwpu.edu.cn
1 Introduction
Multilingual dialogue research has garnered significant attention Xue et al. (2021); Liu et al. (2020); Conneau et al. (2020), especially with the explosive growth of large language models (LLMs) like Touvron et al. (2023); Yang et al. (2024). The development of highly capable multilingual models is urgently needed, which undoubtedly requires high-quality multilingual dialogue datasets. However, while multilingual datasets have been developed for various NLP tasks Lin et al. (2021); Liu et al. (2023); Ponti et al. (2020); Muennighoff et al. (2022), there remains a stark scarcity of such datasets for dialogue generation tasks. The existing dialogue datasets are primarily focused on simple dialogue scenarios and are limited to a few high-resource languages, lacking support for complex multi-party, multi-turn dialogues Hu et al. (2019) and low-resource languages, as shown in Table 1. As a result, there exists a significant gap in the "multilingual + complex dialogue scenarios" domain, which has seriously hindered the development of multilingual dialogue. Furthermore, this gap raises doubts about the adaptability of current LLMs to multilingual settings in complex dialogue scenarios.
Our previous experiments revealed that current LLMs exhibit significant shortcomings in adapting to complex multilingual dialogue scenarios. Specifically, similar to the setting of Choenni et al. (2023), we conducted in-context learning (ICL) experiments using templates and examples in various languages for multi-party dialogue scenarios. We observed that, under various template configurations, increasing the number of examples often led to degraded performance—this was even true for English corpora. The phenomenon indicates that the large models themselves struggle with adapting to MCDS.
Dataset Task Type Sample Count Parallel Corpus Multilingual
UDC RG 320K No No
XCOPA MCQA 7.5k No Yes
XWinograd Reasoning 4.5K No Yes
XDailyDialog RR 52k Yes Yes
Table 1: Common multilingual or multi-party datasets. UDC refers to the Ubuntu Dialogue Corpus. Task types include RG for Response Generation, RR for Response Retrieval, and MCQA for Multiple-Choice Question Answering.
To address this gap, we present XMP, a high-quality parallel multilingual dataset for multi-party dialogue. The dataset is built from real-world multi-speaker interactions in podcasts, and the original English data has been carefully translated into six languages: English, Chinese, Japanese, German, French, and Italian. The dataset consists of 312k conversations spanning diverse topics such as society, culture, politics, and entertainment. To ensure translation quality, we implemented a rigorous quality control process, including systematic data filtering and cleaning to guarantee the diversity and semantic coherence of the dialogue content.
Through extensive experiments, we found that several previously recognized multilingual capabilities Doddapaneni et al. (2021); Wang et al. (2019); Whitehouse et al. (2023) of LLMs no longer hold true on this dataset. For instance, the commonly accepted multilingual complementary ability—where combining English data with another language during fine-tuning significantly boosts the model’s performance on that language—was notably absent in our experiments. Specially, whether working with languages from the same language family (e.g., English + French) or different language families (e.g., English + Chinese), the results on our dataset were inferior to those obtained by training on a single language. Our key contributions are summarized as follows:
  • We construct and release XMP, the first high-quality parallel multilingual dataset for multi-party dialogue, covering six languages and over 312k conversations. The dataset offers diverse topics and supports research into complex multilingual dialogue generation and cross-lingual transfer.
  • We conduct extensive evaluations on XMP, revealing significant limitations in the multilingual capabilities of current LLMs, particularly their inability to leverage multilingual complementary effects in complex dialogue settings.
  • We provide detailed analysis and insights into the mechanisms of LLMs in multilingual environments, highlighting key challenges and opportunities for improving their adaptability to real-world, multi-party dialogue scenarios.
2 Related Work
2.1 Multilingual Datasets for Dialogue
The development of high-quality multilingual dialogue datasets has significantly contributed to the advancement of multilingual research. In recent years, multilingual datasets have expanded to cover a wide range of tasks, such as XCOPA Ponti et al. (2020), which focuses on multiple-choice question answering, and XWinograd Muennighoff et al. (2022), which emphasizes reasoning tasks. However, dialogue-related datasets remain scarce and are often constrained to simple dialogue scenarios. For instance, XDailyDialog Liu et al. (2023) provides a multilingual parallel dialogue dataset with four languages, designed for response retrieval tasks. Similarly, XPersona Lin et al. (2021) extends the Persona-Chat dataset to six languages, focusing on cross-lingual personalized dialogue. While these datasets support tasks such as monolingual and cross-lingual dialogue, they are limited in complexity, often focusing on straightforward conversation patterns and neglecting multi-party and multi-turn dialogue structures. Moreover, they tend to overlook the crucial task of dialogue response generation, which is fundamental for building robust dialogue systems.
2.2 Multilingual Capabilities of Large Language Models
Recent advancements in large language models (LLMs), such as GPT-4 Achiam et al. (2023) and LLaMA Touvron et al. (2023), have showcased impressive multilingual capabilities. By leveraging instruction tuning, human feedback alignment Lai et al. (2024), and vast multilingual training corpora, these models have significantly improved performance across diverse languages, including many low-resource ones. However, their evaluation has predominantly focused on relatively structured tasks, rather than the more dynamic and context-rich settings of dialogue generation.
Multilingual complementarity, a key feature of cross-lingual transfer, has been extensively studied in the context of multilingual language models (MLMs). Choenni et al. (2023) investigated cross-lingual data sharing during fine-tuning and demonstrated that training samples from one language can reinforce or complement the knowledge acquired from other languages. Similarly, Doddapaneni et al. (2021) highlighted that multilingual models often achieve better performance by leveraging linguistic similarities across languages, enabling effective zero-shot and few-shot cross-lingual transfer. These findings suggest that MLMs can effectively exploit shared representations across languages for structured tasks, such as natural language inference and multiple-choice question answering.
However, despite these promising results, the adaptability of large language models (LLMs) to complex dialogue scenarios remains underexplored. Most existing research focuses on relatively structured tasks and overlooks dynamic settings like multi-turn, multi-party dialogue. Our findings indicate that multilingual complementarity often diminishes in these complex dialogue scenarios. For instance, training a model on both English and another language (e.g., French or Chinese) does not consistently outperform single-language training, even for linguistically similar languages.
3 Preliminary Experiments
To investigate the capabilities of large language models (LLMs) in generating responses for complex dialogue scenarios, we conducted preliminary experiments using approximately 1,000 English multi-party dialogue samples, each involving three or more speakers discussing topics such as culture, society, and entertainment. By testing various templates and different numbers of in-context learning (ICL) examples, we observed that LLMs failed to achieve significant improvements from ICL. Even for English, the language LLMs are typically most proficient in, the results showed almost no variation in performance as the number of ICL examples increased, with automatic metrics such as BLEU and ROUGE remaining nearly unchanged.
The phenomenon raised doubts about the ability of LLMs to handle complex dialogue scenarios. What is more, if their performance on English was already so limited, would it be even worse for other languages? To further explore this, we expanded the dataset by translating the dialogues into additional languages. Following the experimental setup in Choenni et al. (2023), we tested the multilingual data under diverse templates and varying ICL example counts. The results revealed a consistent performance decline for all non-English languages as the number of ICL examples increased, regardless of template design. The detailed experimental setup and results are provided in Appendix A.
These findings prompt deeper questions about the nature of LLMs’ capabilities and why ICL fails to yield improvements in this scenario. Much of LLMs’ strength comes from the immense volume of training data they are exposed to, enabling them to approximate conversational patterns and generate contextually plausible responses. However, in complex multi-party dialogue settings, where multiple speakers interact dynamically and linguistic nuances vary significantly across languages, the models’ reliance on data-driven pattern matching seems insufficient. They struggle to incorporate additional examples to refine their understanding of the dialogue’s structure or adapt their reasoning to the specific task context.
4 Data Collection
The Multilingual Complex Dialogue Dataset aims to provide high-quality parallel data for studying multilingual in complex multi-party scenarios. The raw data was collected from podcasts and subsequently translated into parallel corpora using Google Translate. In this section, we describe the process of collecting and preparing the dataset.
4.1 Data Source
In our data collection process, we chose podcasts as the primary data source due to their ability to provide high-quality transcripts and their frequent use of multi-party interaction formats involving three or more speakers. The format aligns well with the needs of multi-party dialogue research and offers a realistic representation of complex dialogue scenarios found in real-world interactions.
We collected data from various publicly available podcasts, including "Life Kit," "Pop Culture Happy Hour," "Short Wave," "Throughline," "Consider This from NPR," "It’s Been a Minute," "The NPR Politics Podcast," "The Lawfare Podcast," and "Planet Money." All these programs are publicly accessible and available for free download. Details regarding the crawling, filtering, cleaning, and segmentation processes are provided in Appendix B.
4.2 Translation Method
After completing the initial collection of English transcripts, the next challenge was expanding them into multiple languages without compromising contextual integrity or the complexity of the dialogue scenarios. The quality of large-scale parallel corpora depends heavily on the accuracy and coherence of the translation at the word, sentence, and conversation levels. To address this, we designed a comprehensive Translation Method Evaluation process to strike an optimal balance between translation quality and efficiency.
Translation Method W-Level U-Level C-Level
Google Translate 92.6 86.9 84.5
Baidu Translate 88.2 73.1 74.7
DeepL Translate 88.7 70.8 73.4
iFLYTEK Translate 86.2 70.2 72.3
LLaMA 3.1 90.4 80.3 82.8
ChatGPT-4o 94.5 85.9 87.8
Human Translation 99.2 98.5 95.4
Table 2: Evaluation Results of Translation Methods
Statistic Train Test Total
#Dialogues 281,616 31,290 312,906
#Utterances 2,384,262 264,888 2,649,150
Avg Turns per Dialogue 8.47 8.48 8.47
Avg Utterance Length 41.88 41.90 41.88
Avg Response Length 55.99 56.33 56.02
Avg Speakers per Dialogue 3.63 3.67 3.64
Avg Turns per Speaker 2.33 2.31 2.33
Table 3: Statistics of our dataset.
We began by selecting 50 representative samples that covered a diverse range of topics (e.g., law, politics, daily life), dialogue turns (ranging from 7 to 15 turns), and multi-party podcast segments. This ensured that the evaluation samples were sufficiently diverse in complexity, providing reliable references for subsequent large-scale translation.
For these samples, we employed several translation methods, including Google Translate, DeepL, Baidu Translate, iFLYTEK Translate, LLaMA 3.1, ChatGPT-4o and human translation. Each method translated the English transcripts into five other languages (Chinese, Japanese, French, German, and Italian), resulting in 300 translations (50 samples × 6 languages). To ensure objectivity and reliability, three linguistics experts reviewed each translation in multiple rounds. The review process identified issues such as lexical misusage, semantic distortion, or contextual inconsistency in the translated outputs. Ultimately, the evaluation focused on three dimensions:
  • Word-level evaluation assessed the accuracy of individual word translations and checked for spelling errors.
  • Utterance-level evaluation examined sentence fluency, semantic accuracy, and appropriateness in context.
  • Conversation-leve evaluation ensured contextual consistency and alignment with the original dataset.
As shown in Table 2, human translation consistently achieved the highest overall quality. However, due to the prohibitive cost and extended timeline of large-scale manual translation, we explored automated tools. Among Google Translate, DeepL, Baidu, iFLYTEK, LLaMA 3, ChatGPt-4o, both ChatGPT-4 and Google Translate have shown impressive results, with ChatGPT-4 slightly outperforming Google Translate. However, ChatGPT-4 requires careful handling of templates, as it may provide natural translations for some examples but sometimes offers literal translations that overlook specific contexts. Considering translation speed and cost-effectiveness, we opted for Google Translate as the more practical choice for large-scale multilingual tasks. Specific case comparisons of the translation tools are detailed in Appendix C.
4.3 Dataset Quality Analysis and Statistics
The dataset includes six languages: English, Chinese, Japanese, French, German, and Italian. While English serves as the original source, all other languages were generated using translation API. To systematically evaluate the quality of the dataset, we conducted a thorough assessment focusing on four key aspects: word-level accuracy, sentence fluency, semantic consistency, and contextual coherence.
To assess the quality of the dataset, we employed expert evaluation. Specially, we randomly selected 500 conversations, covering various topics. For each language, two evaluation experts were selected based on their proficiency in both English and the target language. The experts were trained on the evaluation process, including specific scoring criteria and illustrative examples for each rating level, to ensure consistency and objectivity. In cases where there was significant disagreement between the two experts’ scores, a third expert was consulted for adjudication. The final scores were aggregated and averaged to provide an overall quality score for each language. Detailed evaluation criteria and the scoring results for each language are provided in Appendix D.
Figure 1: Distribution of Speaker Counts
In Table 3, we present the detailed statistical analysis of the dataset. The dataset contains a total of 312k dialogues and 2,649k utterances, with an average of 8.47 turns per dialogue, indicating a high level of dialogue density. This ensures that each dialogue typically involves multiple exchanges, creating scenarios rich in multi-turn interactions. The dataset also emphasizes its multi-party characteristics, with an average of 3.64 speakers per dialogue. Notably, over 75% of the dialogues involve at least three speakers, as shown in Figure 1, reflecting the interaction between multiple distinct roles. This multi-party nature enhances the realism of the dataset and provides a robust foundation for studying complex dialogue scenarios. Moreover, the average number of turns per speaker is 2.33, indicating that most speakers contribute multiple times to the conversation. Additionally, we ensured that the target speaker (i.e., the speaker for whom a response is to be generated) had spoken at least once earlier in the dialogue history. This design guarantees that the model has sufficient contextual information to generate coherent and relevant responses.
5 Experiments
5.1 Dataset Split and Baseline Models
For training and evaluating multilingual complex dialogue scenarios, we utilized the XMP dataset, splitting it into 90% training and 10% testing. The dataset includes six languages: English, Chinese, Japanese, French, German, and Italian, covering a wide range of topics and multi-party interactions, which provides a challenging benchmark for multilingual dialogue generation tasks.
To benchmark the performance of state-of-the-art large language models (LLMs) on the XMP dataset, we selected three open-source instruction-tuned models as baselines: LLaMA 3.1 8B Instruct, Qwen 2.5 7B Instruct, and LLaMA 3.1 70B Instruct. These models were chosen for their demonstrated capabilities in multilingual dialogue response task. Their inclusion allows us to examine the trade-offs between model size, and response quality across six languages
5.2 Task Setup
The generation task focuses on producing high-quality responses in complex multi-party dialogue scenarios. Given a conversational context and the current speaker, the model is tasked with generating a coherent, contextually relevant, and linguistically accurate response.
Specially, Let $D={(H,S,R)}^{n}_{i=1}$represent the multilingual dialogue dataset for complex scenarios, where $H$ is the dialogue history composed of multiple “Speaker: Utterance” pairs concatenated sequentially, $S$ is the current speaker, and $R$ is the ground truth response. To guide the model in generating speaker-specific responses, a prompt is dynamically constructed based on the current speaker, such as “What will Speaker-x say next?”. The dialogue history is then concatenated with this prompt to form the input for the model. During training, the model predicts the response $\hat{R}$, and the cross-entropy loss is computed between and the ground truth $L=-\frac{1}{m}{\textstyle\sum_{j=1}^{m}}R_{j}log\hat{R}_{j}$, where $m$ is the number of tokens in $R$.
5.3 Implementation Details
We utilized LLaMAFactory as the training framework and applied LoRA (Low-Rank Adaptation) for efficient fine-tuning. The training was conducted on NVIDIA A100 GPUs using the AdamW optimizer with a learning rate of $1e^{-5}$, a warm-up of 500 steps, and a batch size of 8. To enhance generalization, we standardized key entities in the dialogue history and replaced real entity data with domain-specific slot placeholders. For inference, we used vLLM, an optimized framework that ensures fast and scalable generation. Detailed training and inference commands, along with parameter settings, are provided in Appendix E.
6 Result and Analysis
Language Model F1 BLEU-1 BLEU-2 BLEU-3 BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L
English (en) Qwen 7B 20.34 -0.89 13.31 -0.11 6.07 -0.04 3.74 -0.12 2.49 -0.13 21.68 -0.91 4.00 -1.60 16.09 -1.65
LLaMA 8B 19.92 -1.31 12.80 -0.62 5.73 -0.38 3.45 -0.41 2.25 -0.37 21.26 -1.33 3.73 -1.87 15.54 -2.20
LLaMA 70B 21.23 13.42 6.11 3.86 2.62 22.59 5.60 17.74
Chinese (zh) Qwen 7B 18.65 -0.49 11.45 -0.22 4.72 -0.29 2.78 -0.63 1.77 -0.89 19.53 -0.84 2.60 -1.45 15.74 -1.15
LLaMA 8B 18.37 -0.77 11.60 +0.23 4.68 -0.33 2.76 -0.65 1.76 -0.90 19.26 -1.11 2.45 -1.60 15.44 -1.45
LLaMA 70B 19.14 11.37 5.01 3.41 2.66 20.37 4.05 16.89
Japanese (ja) Qwen 7B 28.17 +1.07 17.74 +1.16 9.27 +0.31 5.98 +0.59 3.96 +0.19 34.76 +1.81 9.16 +0.32 19.08 -0.99
LLaMA 8B 27.69 +0.59 16.81 +0.23 8.76 -0.20 5.72 -0.67 3.90 -0.87 33.71 +0.76 8.73 -1.11 19.08
LLaMA 70B 27.10 15.58 8.96 6.39 4.77 32.95 9.84 20.07
French (fr) Qwen 7B 16.35 -1.13 10.14 -0.27 4.31 -0.09 2.59 -0.25 1.68 -0.13 17.44 -1.25 2.41 -1.63 13.20 -1.78
LLaMA 8B 17.20 -0.28 11.21 +0.80 4.87 +0.47 3.00 +0.16 2.00 +0.19 18.34 -0.35 2.85 -1.19 13.63 -1.35
LLaMA 70B 17.48 10.41 4.40 2.84 1.81 18.69 4.04 14.98
German (de) Qwen 7B 17.36 -1.09 11.06 -0.34 4.68 -1.03 2.76 -1.19 1.77 -1.22 17.89 -1.23 2.55 -1.48 14.21 -1.72
LLaMA 8B 18.09 -0.36 12.05 +0.65 5.18 -0.53 3.07 -0.88 2.01 -0.98 18.54 -0.58 2.86 -1.17 14.53 -1.40
LLaMA 70B 18.45 11.40 5.71 3.95 2.99 19.12 4.03 15.93
Italian (it) Qwen 7B 16.42 -1.01 10.26 -0.41 4.20 -1.27 2.49 -1.40 1.59 -1.40 17.03 -1.13 2.09 -1.75 13.38 -1.64
LLaMA 8B 17.12 -0.31 11.01 +0.34 4.65 -0.82 2.81 -1.08 1.83 -1.16 17.70 -0.46 2.48 -1.36 13.79 -1.23
LLaMA 70B 17.43 10.67 5.47 3.89 2.99 18.16 3.84 15.02
Table 4: Performance comparison of Qwen-2.5 7B, LLaMA-3.1 8B, and LLaMA-3.1 70B across different languages. The differences between Qwen-2.5 7B and LLaMA-3.1 8B relative to LLaMA-3.1 70B for each metric are calculated and highlighted.
6.1 Model Performance
As shown in Table 4, we evaluate Qwen-2.5 7B, LLaMA-3.1 8B, and LLaMA-3.1 70B on the XMP dataset across six languages: English, Chinese, Japanese, French, German, and Italian. For text tokenization, Jieba was used for Chinese, Janome for Japanese, and NLTK for the remaining languages.
English outperforms other languages on most metrics, except for Japanese. The result aligns with expectations, as English occupies a significant proportion of the pretraining data for these models. The results highlight the over-reliance of LLMs on high-resource languages.
The performance gain of LLaMA-3.1 70B over smaller models (e.g., 8B or Qwen-2.5 7B) is marginal, with improvements typically within 1% across most metrics. This suggests that the emergent abilities often associated with larger models are not prominently reflected in this task. Instead, the slight advantage may stem from the larger models’ capacity to generate grammatically and lexically refined text, rather than demonstrating a better understanding of the task.
Additionally, Japanese yields unexpectedly high scores across all models. Despite its low proportion in pretraining corpora, Japanese outperforms other non-English languages. We hypothesize that this stems from the unique syntactic structure and word composition of Japanese, which relies heavily on fixed sentence patterns and the use of kana. Tokenization using Janome likely also contributes to higher n-gram overlap, boosting surface-level metrics like BLEU and ROUGE. To validate this hypothesis, we included BERTScore as an additional evaluation metric. Unlike n-gram metrics, BERTScore evaluates semantic similarity.
6.2 Semantic Analysis with BERTScore
Language BERTScore
EN 0.8507
ZH 0.6086
JA 0.6730
FR 0.6707
DE 0.6640
IT 0.6657
Table 5: BERTScore results for LLaMA-3.1 8B across different languages.
Table 4 presents BERTScore results for all six languages. English achieves the highest semantic alignment, confirming its strong representation in training data. Japanese, despite its high BLEU and ROUGE scores, shows comparable BERTScore results to other non-English languages, indicating that its apparent advantage in traditional metrics stems from tokenization effects rather than true semantic superiority.
For other languages, the results indicate that Japanese achieves a comparable BERTScore to other non-English languages, suggesting that its seemingly superior performance in traditional metrics reflects tokenization effects rather than genuine semantic alignment. In contrast, Chinese scores similarly to other non-English languages in metrics such as F1, BLEU, ROUGE-1, ROUGE-2, and ROUGE-L, yet records the lowest BERTScore at just 0.6086. This discrepancy indicates significant differences in structural and semantic expression between Chinese and other languages. Unlike many other languages, Chinese has less rigid grammatical requirements, allowing sentences with entirely different meanings to use similar words and word orders.
6.3 Multilingual Complementary Ability
Train Data Test Data F1 BLEU-1 BLEU-2 BLEU-3 BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L
LLaMA-8B
DE (single) DE 18.09 12.05 5.18 3.07 2.01 18.54 2.86 14.53
DE, FR (mixed) DE 16.58-1.51 9.87-2.18 4.25-0.93 2.64-0.43 1.79-0.22 17.34-1.20 2.40-0.46 14.06-0.47
DE, ZH (mixed) DE 16.49-1.60 9.72-2.33 4.15-1.03 2.56-0.51 1.72-0.29 17.28-1.26 2.36-0.50 13.94-0.59
EN, DE (mixed) DE 16.48-1.61 9.86-2.19 4.14-1.04 2.52-0.55 1.66-0.35 17.24-1.30 2.23-0.63 13.94-0.59
FR (single) FR 17.20 11.21 4.87 3.00 2.00 18.34 2.85 13.63
DE, FR (mixed) FR 15.66-1.54 9.15-2.06 3.99-0.88 2.53-0.47 1.71-0.29 16.79-1.55 2.34-0.51 13.19-0.44
EN, FR (mixed) FR 15.59-1.61 9.02-2.19 3.94-0.93 2.48-0.52 1.66-0.34 16.73-1.61 2.33-0.52 13.10-0.53
ZH (single) ZH 18.37 11.60 4.68 2.76 1.76 19.26 2.45 15.44
DE, ZH (mixed) ZH 17.70-0.67 10.24-1.36 4.29-0.39 2.65-0.11 1.76+0.00 18.89-0.37 2.43-0.02 15.45+0.01
EN, ZH (mixed) ZH 17.69-0.68 10.18-1.42 4.23-0.45 2.57-0.19 1.68-0.08 18.90-0.36 2.37-0.08 15.36-0.08
EN (single) EN 19.92 12.80 5.73 3.45 2.25 21.26 3.73 15.54
EN, DE (mixed) EN 19.73-0.19 12.25-0.55 5.79+0.06 3.73+0.28 2.62+0.37 21.17-0.09 4.08+0.35 16.13+0.59
EN, FR (mixed) EN 19.74-0.18 12.26-0.54 5.77+0.04 3.68+0.23 2.54+0.29 21.18-0.08 4.10+0.37 16.14+0.60
EN, ZH (mixed) EN 19.70-0.22 12.18-0.62 5.74+0.01 3.70+0.25 2.57+0.32 21.27+0.01 4.10+0.37 16.12+0.58
LLaMA-70B
DE (single) DE 18.45 11.40 5.71 3.95 2.99 19.12 4.03 15.93
DE, FR (mixed) DE 18.45-0.00 11.44+0.04 5.70-0.01 3.85-0.10 3.05+0.06 19.58+0.46 4.30+0.27 16.23+0.30
DE, ZH (mixed) DE 18.49+0.04 11.37-0.03 5.80+0.09 3.91-0.04 3.06+0.07 19.01-0.11 4.15+0.12 15.88-0.05
EN, DE (mixed) DE 17.00-1.45 11.70+0.30 5.12-0.59 3.08-0.87 2.06-0.93 17.79-1.33 2.97-1.06 13.30-2.63
FR (single) FR 17.48 10.41 5.40 3.84 2.96 18.69 4.04 14.98
DE, FR (mixed) FR 18.14+0.66 11.20+0.79 5.88+0.48 4.16+0.32 3.17+0.21 19.26+0.57 4.45+0.41 15.40+0.42
EN, FR (mixed) FR 18.59+1.11 12.14+1.73 5.98+0.58 4.04+0.20 2.99+0.03 19.42+0.73 4.11+0.07 15.18+0.20
ZH (single) ZH 19.14 11.37 5.70 4.01 3.06 20.37 4.05 16.89
DE, ZH (mixed) ZH 19.06-0.08 11.33-0.04 5.74+0.04 4.10+0.09 3.12+0.06 20.33-0.04 4.25+0.20 17.17+0.28
EN, ZH (mixed) ZH 19.08-0.06 11.26-0.11 5.53-0.17 4.01-0.00 3.03-0.03 20.30-0.07 4.13+0.08 17.07+0.18
EN (single) EN 21.23 13.42 7.11 4.98 3.76 22.59 5.60 17.74
EN, DE (mixed) EN 17.83-3.40 11.87-1.55 4.90-2.21 2.81-2.17 1.76-2.00 19.37-3.22 2.77-2.83 13.10-4.64
EN, FR (mixed) EN 21.94+0.71 14.52+1.10 7.69+0.58 5.40+0.42 4.10+0.34 23.17+0.58 5.91+0.31 17.98+0.24
EN, ZH (mixed) EN 21.67+0.44 13.51+0.09 7.18+0.07 4.96-0.02 3.72-0.04 23.27+0.68 5.61+0.01 17.76+0.02
Table 6: Comparison of multilingual complementary ability for LLaMA-3.1 8B and LLaMA-3.1 70B, grouped by testing language. Results include single- and mixed-language training scenarios.
Multilingual complementary ability is defined as a model’s capacity to effectively leverage information from multiple languages during training to enhance performance on specific target languages. The results of our experiments, as detailed in Table 6, shed light on this phenomenon by using LLaMA-3.1 8B and LLaMA-3.1 70B models tested on the XMP dataset. Notably, our findings diverge significantly from conventional understanding in prior research.
6.3.1 Mixed Training with English
Previous research often highlights that mixing English as a source and target language during training provides significant benefits for the target language due to English’s dominance in pretraining corpora and its rich semantic structure. However, our experiments on XMP challenge this assumption. For the LLaMA-3.1 8B model, mixing English with any other language consistently leads to negative performance impacts. Interestingly, for the LLaMA-3.1 70B model, this detrimental effect is mitigated, particularly when English is mixed with French. However, even in cases where certain languages exhibit improvements in specific metrics, these gains remain minimal, typically less than 0.5%. This suggests that while larger model capacities can temper the negative impacts of mixed training, the benefits derived from cross-lingual interaction are marginal and insufficient to fully leverage the potential of multilingual complementary ability in complex dialogue scenarios.
6.3.2 Cross-Language Benefits Are Asymmetrical
The experimental results for 70B-model reveal an intriguing asymmetry: while English typically fails to enhance the performance of other languages, other languages, such as French or German, often provide improvements when mixed with English. It might stem from the model’s reliance on English during pretraining, making it more adept at leveraging additional multilingual context for English-centric tasks while struggling to adapt English’s resources to enhance other languages effectively.
6.3.3 Effect of Linguistic Similarity
The relationship between linguistic similarity and cross-lingual gains appears to be more complex than previously assumed. Traditionally, languages within the same linguistic family, such as English and German, or French and English, are expected to exhibit significant cross-lingual transfer benefits when trained together. However, our experimental findings suggest that these linguistic connections do not guarantee consistent or significant cross-lingual improvements. For instance, the mixing of English and German often results in negative impacts on both languages’ performance, even though they are closely related in terms of linguistic roots. Conversely, the pairing of French with either German or English generally produces positive outcomes, albeit with varying degrees of improvement depending on the specific metrics and target languages.
The nuanced relationship highlights the need for deeper investigations into how linguistic similarity interacts with other variables in multilingual training, particularly in complex and dynamic tasks. It underscores that shared linguistic heritage, while beneficial, is not a deterministic factor for cross-lingual success, and the benefits of multilingual training are far from uniform across different language pairs
7 Conclusion
In this study, we introduced XMP, a high-quality multilingual parallel multi-party dialogue dataset, designed to address the gaps in existing datasets for complex dialogue scenarios. XMP provides extensive coverage of six languages to address the lack of comprehensive resources for complex multilingual dialogue generation. Using XMP as a benchmark, we evaluated the multilingual complementary ability of LLaMA-3.1 8B and 70B models, uncovering unexpected findings that challenge prior research. Mixing English with other languages during training often resulted in performance degradation, particularly for smaller models, while larger models like the 70B variant demonstrated better resilience and achieved marginal gains in specific cases, such as English-French mixing. Interestingly, linguistic similarity did not always yield mutual reinforcement; for instance, English and German showed negative impacts when mixed, whereas French consistently enhanced performance when paired with other languages. These insights underscore the limitations of current multilingual models and highlight the need for further research into linguistic properties and cross-lingual transfer, with XMP serving as a robust benchmark for advancing multilingual dialogue generation.
8 Limitations
Despite the contributions of this study, several limitations warrant discussion. First, the XMP dataset currently includes a limited number of languages, omitting many low-resource languages, which constrains its generalizability to broader multilingual contexts. Second, during the dataset collection process, the team also collected corresponding podcast audio data. However, the significant challenges associated with audio preprocessing have so far prevented XMP from being expanded into a multimodal corpus. This limitation is a priority for future improvement, with plans to incorporate audio features to enhance the dataset’s utility. Finally, the exploration of linguistic similarity in this study was limited to a few language pairs, primarily focusing on German, French, English, and Chinese, leaving the effects of cross-lingual interactions among other linguistic families underexplored. Addressing these limitations could significantly enhance the applicability and impact of XMP in future multilingual and multimodal research.
Appendix A Preliminary Experiments
Before constructing the full dataset, we first conducted a small-scale collection of approximately 1,000 English multi-party dialogue samples. Each sample contained at least three speakers and covered a variety of topics to ensure diversity and complexity in the dialogues. To evaluate the large model’s multi-party dialogue response generation capability, we used In-Context Learning (ICL) testing. We found that as the number of examples in the ICL increased, there was virtually no improvement in the quality of the model’s generated responses. This was particularly evident with the English data, which led us to question how the large model would perform on relatively sparse, under-pretrained corpora. To investigate this further, we expanded the 1,000 English samples to include French, German, Italian, and Chinese.
Figure 2: This is an case with the target language in Chinese, showcasing three different template types (Type 1, Type 2, Type 3).
Figure 3: The results of Preliminary Experiments. $\{metric\}\_{n}$ means n examples used in the ICL and $\{Language\}\_{x}$ means Type X template used.
Following the approach of CCD et al., we applied different templates combined with ICL for further validation. Specially, we designed three different experimental methods to validate the model’s performance variations:
  • 1.
    description Using English templates and English examples. In this setting, the model learns and infers based on English examples, effectively leveraging the provided data.
  • 2.
    Using target language templates and target language examples. For example, when testing with Chinese data, we employ Chinese templates and Chinese examples.
  • 3.
    Using English templates and target language examples.
We also experimented with various template formats, but the results were largely similar across different configurations. As a result, we selected one template format that was used in the final experiments, which is illustrated in the Figure 2. Additionally, we tested the model’s performance with different numbers of examples (0, 1, 2, 4, 8).
The results are shown in Figure 3. Regardless of the template configurations used, we observed that as the number of examples in the ICL increased, the evaluation metrics generally decreased. These results raise important questions about the capabilities of LLMs and why ICL fails to lead to improvements in this context. LLMs are typically strong due to the vast amount of data they are trained on, which helps them learn conversational patterns and produce contextually appropriate responses. However, in the complex environment of multi-party dialogues, where several speakers interact in dynamic ways, the models’ dependency on data-driven pattern matching falls short. They face challenges in effectively incorporating additional examples to enhance their understanding of the dialogue structure or adjust their reasoning to the specific context of the task.
Appendix B Data Collection and Preprocessing
This chapter details the process of data collection, filtering, cleaning, and segmentation to construct our high-quality multilingual dataset.
B.1 Data Collection
We selected podcasts as our primary data source due to their ability to provide authentic conversational content. Most podcast episodes involve discussions among three or more participants, effectively simulating real-world multi-party interaction scenarios. Data was collected from publicly accessible podcast web: www.npr.org, including Life Kit, Pop Culture Happy Hour, Short Wave, Throughline, Consider This from NPR, It’s Been a Minute, The NPR Politics Podcast, The Lawfare Podcast, and Planet Money. All these programs are in the public domain, freely available for access and download.
The complete transcription content of the podcasts was fetched by constructing and sending HTTP requests to the relevant pages. HTML responses were parsed to extract transcription data. Specifically:
  • 1.
    The HTML structure was parsed to extract the page titles.
  • 2.
    The <div> sections containing transcription text were identified and processed.
  • 3.
    Extraneous content, such as timestamps and bracketed text, was removed or converted into readable text.
  • 4.
[…]
    was applied to separate "Speaker: Utterance" pairs, aligning speakers with their corresponding utterances.
Type Translated script
Origin An untried producer and his cast of complete unknowns prepare to make television history while network suits breathe down their necks. And just about everything that can go wrong does, like, completely, totally. It’s bad. I’m Aisha Harris.
Google {CJK}UTF8gbsn一位未经验考验的制片人和他的一众无名小卒准备创造电视历史,而电视网络高层对他们紧追不舍。几乎所有可能出错的事情都完全出错了。这很糟糕。我是艾莎·哈里斯。
Baidu {CJK}UTF8gbsn一位未经验考验的制片人和他的一群完全未知的人准备创造电视历史,而网络服则压在他们的脖子上。几乎所有可能出错的事情会完全、彻底地发生。这很糟糕。我是艾莎·哈里斯。
DeepL {CJK}UTF8gbsn一位未经验考验的制片人和他的一帮完全不知名的演员准备创造电视历史,而电视台的西装革履让他们喘不过气来。一切可能出错的地方都出了错,就像,完全,彻底。这很糟糕。我是艾莎·哈里斯。
iFlytek {CJK}UTF8gbsn一位未经验考验的制片人和他的一帮完全不知名的演员准备创造电视历史,而电视台的西装束紧着他们。几乎所有可能出错的事情会发生,就像,完全,彻底。这很糟糕。我是艾莎·哈里斯。
LLaMA3 {CJK}UTF8gbsn一位毫无经验的制片人和他全新演员阵容准备创造电视史,可电视台的高层却紧盯着他们。几乎所有可能出错的地方都出了问题,真是糟透了。我是艾莎·哈里斯。
Human {CJK}UTF8gbsn一位未尝试的制片人带领一群毫无名气的演员,准备改写电视规则,可电视台高层却紧盯着他们的一举一动。结果几乎出错的地方都有错,真是糟糕了。我是艾莎·哈里斯。
Table 7: Specific case comparisons of the translation tools.
B.2 Data Filtering
Post extraction, a set of filtering rules was applied to eliminate noise from the data:
  • 1.
    Platform Labels: Phrases such as “Copyright,” “(Laughter),” “(SOUNDBITE OF TV SHOW, THE TONIGHT SHOW),” and “HOST, BYLINE” were removed.
  • 2.
    URLs: Any URL strings within the text were removed.
B.3 Data Segmentation
Podcast episodes typically range from 80 to 100 utterances, making segmentation essential to generate manageable samples. The following rules were used:
  • 1.
    Segment Length: Long conversations were split into shorter segments, ensuring each contained 7 to 15 utterances.
  • 2.
    Speaker Presence: The speaker generating the response must appear in the prior dialogue history to ensure coherence.
  • 3.
    Response Quality: Segments with overly short or long responses, or responses deemed meaningless (e.g., “Yeah, wel” or “Alright”), were excluded.
Appendix C Chosen of Translate API
Origin Prompt 1:{CJK}UTF8gbsn翻译成中文 Prompt 2:{CJK}UTF8gbsn翻译成中文,请不要直译,在这句话中有很多需要理解并意译的地方,请注意
An untried producer and his cast of complete unknowns prepare to make television history while network suits breathe down their necks. And just about everything that can go wrong does, like, completely, totally. It’s bad. I’m Aisha Harris. {CJK}UTF8gbsn一位未经试验的制片人和他的完全不为人知的演员阵容准备创造电视历史,而电视网络的高层正对他们施加巨大压力。几乎所有可能出错的事情都出了错,完全、彻底地错了。情况非常糟糕。我是Aisha Harris。 {CJK}UTF8gbsn一位经验不足的制片人和一群完全没有名气的演员,正准备创造电视历史,而电视台的高层则在背后施压。几乎所有可能出错的事都发生了,而且是彻底的、完全的错误,局面一团糟。我是Aisha Harris。
Table 8: This is the translation result by ChatGPT-4o. As you can see, using different prompts leads to completely different translations. The result with prompt 2 is much better than result with prompt 1.
To further illustrate our reasoning for choosing the Google Translate API, the Table 7 and Table C present an actual translation snippet of a dialogue. In terms of word-level evaluation, all the translation APIs demonstrated accurate spelling without errors. Google’s translation of "cast of complete unknowns" as "{CJK}UTF8gbsn 一众无名小卒" is both vivid and precise. In contrast, Baidu’s "{CJK}UTF8gbsn网络服装" is a literal translation of "network suits," failing to capture the intended meaning. Similarly, DeepL’s "{CJK}UTF8gbsn西装革履" and iFlytek’s "{CJK}UTF8gbsn西装" translations of "network suits" are overly literal, lacking the semantic nuance of referring to "network executives."
For utterance-level evaluation, Google’s rendering of "{CJK}UTF8gbsn电视网络高层对他们紧追不舍" effectively conveys the urgency in "network suits breathe down their necks," while maintaining fluency and accuracy in alignment with the original context. Baidu’s translation of "{CJK}UTF8gbsn一位未经尝试的制片人" is less precise and deviates from the original meaning. The phrase "{CJK}UTF8gbsn完全、彻底地发生" is grammatically correct but verbose and unnatural. Overall, Baidu’s translation style tends to be rigid, lacking smooth and natural expression. DeepL and iFlytek’s translations, such as "{CJK}UTF8gbsn就像,完全,彻底," fail to accurately reflect the overall semantics and lack fluency. Meanwhile, LLaMA3’s translation of "cast of complete unknowns" as "{CJK}UTF8gbsn全新演员阵容" shifts the original "unknown" connotation to a more positive tone, which is inconsistent with the original context.
For conversation-level evaluation, Google Translate demonstrates coherence and effectively conveys the original meaning with high parallelism, maintaining a natural conversational tone. While LLaMA3 generally performs well in contextual consistency, certain examples, such as this one, show occasional mismatches in contextual understanding.
Comprehensive Evaluation: The Google Translate API outperforms other tools and models by better capturing the urgency and fluency of the original text. Although there are still slight deficiencies compared to human translation, it is sufficiently effective for large-scale multilingual parallel data construction. While LLaMA3 translations also exhibit reasonable quality, they sometimes produce contextually irrelevant outputs or fail to handle languages with scarce training data effectively, generating extraneous or unfaithful translations.
Appendix D Dataset Quality: Detailed evaluation criteria and the scoring results
To assess the quality of the dataset, we employed expert evaluation. Specifically, we randomly selected 500 conversations, covering various topics. The review was conducted based on the following aspects, with scores ranging from 1 to 5 for each metric. The final scores and detailed scoring criteria are shown in Table 9.
Metric Score Range Description Final Score
Conversation Diversity 1 Very Low Diversity: Narrow range of topics, repetitive dialogue structures, limited interaction patterns. 4.74
2 Low Diversity: Few different topics, minimal variation in interaction styles, some repetition.
3 Moderate Diversity: A reasonable range of topics, some diversity in interaction styles and contexts.
4 High Diversity: Diverse topics and contexts, multiple types of dialogues and varied interactions.
5 Very High Diversity: Rich variety of topics, multiple communication styles, diverse contexts and subjects.
Conversation Fluency 1 Very Low Fluency: Disjointed conversations with abrupt topic changes, frequent grammatical errors. 4.21
2 Low Fluency: Awkward transitions, slight disruptions in flow, minor grammatical issues.
3 Moderate Fluency: Clear dialogue with occasional awkward phrasing, understandable overall.
4 High Fluency: Natural flow, clear transitions between topics, minor errors not affecting overall fluency.
5 Very High Fluency: Seamless conversation with smooth transitions, grammatically correct throughout.
Annotation Quality 1 Very Low Quality: Responses are either too short or excessively long, often irrelevant or nonsensical. Responses frequently deviate from the previous context, making them incoherent or disjointed. 4.43
2 Low Quality: Responses are occasionally too short or too long, with some parts being irrelevant. There are noticeable inconsistencies with the previous context, leading to slight confusion in the conversation flow.
3 Moderate Quality: Responses are generally of appropriate length but may occasionally lack relevance or slightly stray from the context. They remain understandable and contextually coherent, with some minor lapses.
4 High Quality: Responses are of reasonable length, relevant, and mostly meaningful. There is good consistency with the previous dialogue, but small deviations or slight misunderstandings may still occur.
5 Excellent Quality: Responses are always appropriately lengthened, coherent, and directly related to the previous dialogue. They are contextually meaningful, contributing effectively to the conversation without unnecessary repetition or confusion.
Table 9: Detailed scoring criteria for dataset quality evaluation.
Appendix E Detailed training and inference command
  • Training qwen-2.5 7b:
[…]

### v2

@@
-Can xLLMs Understand the Structure of Dialog? Exploring Multilingual Response Generation in Complex Scenarios
+Can MLLMs Generalize to Multi-Party dialog? Exploring Multilingual Response Generation in Complex Scenarios
 Abstract
-Multilingual research has garnered increasing attention, especially in the domain of dialogue systems.
-The rapid advancements in large language models (LLMs) have fueled the demand for high-performing multilingual models.
-However, two major challenges persist: the scarcity of high-quality multilingual datasets and the limited complexity of existing datasets in capturing realistic dialogue scenarios.
-To address these gaps, we introduce XMP, a high-quality parallel Multilingual(X) dataset sourced from Multi-party Podcast dialogues [^1] https://github.com/ahxchzt/XMP_data.
-Each sample in the dataset features at least three participants discussing a wide range of topics, including society, culture, politics, and entertainment.
-Through extensive experiments, we uncover significant limitations in previously recognized multilingual capabilities of LLMs when applied to such complex dialogue scenarios.
-For instance, the widely accepted multilingual complementary ability of LLMs is notably impacted.
-By conducting further experiments, we explore the mechanisms of LLMs in multilingual environments from multiple perspectives, shedding new light on their performance in real-world, diverse conversational contexts.
-Can xLLMs Understand the Structure of Dialog? Exploring Multilingual Response Generation in Complex Scenarios
+Current multilingual large language models(MLLMs) still focus on simple question-answering formats, often overlooking more complex dialogue scenarios.
+In other words, their capabilities of multilingual large models have yet to be validated in dialogue tasks with intricate structures.
+We therefore ask, Q1: How well do LLMs generalize to more complex dialog scenarios?
+Q2: Can supervised fine-tuning on a high-quality parallel benchmark restore this ability?
+Q3: Does the "multilingual complementarity" effect survive in the setting?
+To answer these questions, we introduce XMP [^1] https://anonymous.4open.science/r/XMP_data-5CC7, a high-quality parallel Multilingual(X) dataset sourced from Multi-party Podcast dialogues, which is the first parallel dataset focusing on multi-party dialogue scenarios.
+Most samples in the dataset feature three or more participants, discussing a wide range of topics.
+Through extensive experiments, we find that, R1: MLLMs fail to generalize to multi-party setting, R2: Fine-tuning on XMP improves only marginally, with the 70B model achieving at most a 1% absolute gain over its 8B counterpart; R3: Mixing languages during SFT is usually detrimental, with any benefits being marginal and limited to isolated cases in the 70B model.
+Can MLLMs Generalize to Multi-Party dialog? Exploring Multilingual Response Generation in Complex Scenarios
 Zhongtian Hu1, Yiwen Cui1, Ronghan Li2, Meng Zhao3, Lifang Wang1, 1School of Computer Science and Engineering, Northwestern Polytechnical University, 2School of Computer Science and Technology, Xidian University, 3School of Artificial Intelligence and Big Data, Henan University of Technology ahxchzt@mail.nwpu.edu.cn, wanglf@nwpu.edu.cn
 1 Introduction
-Multilingual dialogue research has garnered significant attention Xue et al.
+Amid the rising demand for seamless cross-lingual communication and the growing complexity of global language needs, multilingual research has garnered widespread attention Xue et al.
 (2021); Liu et al.
 (2020); Conneau et al.
-(2020), especially with the explosive growth of large language models (LLMs) like Touvron et al.
+(2020a).
+Expectations for large language models (LLMs) have grown substantially, Touvron et al.
 (2023); Yang et al.
-(2024).
-The development of highly capable multilingual models is urgently needed, which undoubtedly requires high-quality multilingual dialogue datasets.
-However, while multilingual datasets have been developed for various NLP tasks Lin et al.
+(2024) particularly their ability to handle a wide range of multilingual NLP tasks Lin et al.
 (2021); Liu et al.
@@
 (2020); Muennighoff et al.
-(2022), there remains a stark scarcity of such datasets for dialogue generation tasks.
-The existing dialogue datasets are primarily focused on simple dialogue scenarios and are limited to a few high-resource languages, lacking support for complex multi-party, multi-turn dialogues Hu et al.
-(2019) and low-resource languages, as shown in Table 1.
-As a result, there exists a significant gap in the "multilingual + complex dialogue scenarios" domain, which has seriously hindered the development of multilingual dialogue.
-Furthermore, this gap raises doubts about the adaptability of current LLMs to multilingual settings in complex dialogue scenarios.
-Our previous experiments revealed that current LLMs exhibit significant shortcomings in adapting to complex multilingual dialogue scenarios.
-Specifically, similar to the setting of Choenni et al.
-(2023), we conducted in-context learning (ICL) experiments using templates and examples in various languages for multi-party dialogue scenarios.
-We observed that, under various template configurations, increasing the number of examples often led to degraded performance—this was even true for English corpora.
-The phenomenon indicates that the large models themselves struggle with adapting to MCDS.
-Dataset Task Type Sample Count Parallel Corpus Multilingual
-UDC RG 320K No No
-XCOPA MCQA 7.5k No Yes
-XWinograd Reasoning 4.5K No Yes
-XDailyDialog RR 52k Yes Yes
-Table 1: Common multilingual or multi-party datasets.
-UDC refers to the Ubuntu Dialogue Corpus.
-Task types include RG for Response Generation, RR for Response Retrieval, and MCQA for Multiple-Choice Question Answering.
-To address this gap, we present XMP, a high-quality parallel multilingual dataset for multi-party dialogue.
-The dataset is built from real-world multi-speaker interactions in podcasts, and the original English data has been carefully translated into six languages: English, Chinese, Japanese, German, French, and Italian.
+(2022).
+Among these tasks, dialogue generation stands out as one of the most crucial.
+However, current LLMs have only demonstrated their multilingual capabilities in relatively simple, two-party dialogue scenarios.
+More complex interaction settings, such as multinational meetings or multi-player role-playing games are largely unexplored.
+Figure 1: An example of a multi-party dialogue illustrating complex conversational dependencies, showcasing turn-taking, contextual references, and opinion shifts.
+To date, no multilingual dataset exists that focuses specifically on more complex dialogue settings involving long contexts and multiple speakers Hu et al.
+(2019).
+An example of such a scenario is illustrated in Figure 1, where the structure and dynamics of the conversation highlight its complexity.
+This absence of suitable datasets creates a critical research gap, severely limiting our understanding of how effectively existing MLLMs generalize to realistic and complex dialogue scenario.
+We systematically address the following clearly defined research questions, Q1: How well do LLMs generalize to more complex dialog scenarios?
+Q2: Can supervised fine-tuning on a high-quality parallel benchmark restore this ability?
+Q3: Prior studies like Mulcaire et al.
+(2019); Conneau et al.
+(2020b); Whitehouse et al.
+(2023) have shown that mixing multiple languages during training can enhance model performance.
+Does this effect still hold under the more complex dialogue settings considered in our work?
+To investigate these issues, we present XMP, a high-quality parallel multilingual dataset for multi-party dialogue.
+The dataset is built from real-world multi-speaker interactions in podcasts, and the original English data has been carefully translated into five languages: Chinese, Japanese, German, French, and Italian.
 The dataset consists of 312k conversations spanning diverse topics such as society, culture, politics, and entertainment.
-To ensure translation quality, we implemented a rigorous quality control process, including systematic data filtering and cleaning to guarantee the diversity and semantic coherence of the dialogue content.
+To ensure translation quality, we implemented a rigorous quality control process, including systematic data filtering and cleaning to maintain the diversity and semantic coherence of the dialogue content.
+In addition, we conducted a comprehensive quality assessment to ensure both cross-lingual consistency and cultural fidelity across all translated dialogues.
 Through extensive experiments, we found that several previously recognized multilingual capabilities Doddapaneni et al.
@@
 (2023) of LLMs no longer hold true on this dataset.
-For instance, the commonly accepted multilingual complementary ability—where combining English data with another language during fine-tuning significantly boosts the model’s performance on that language—was notably absent in our experiments.
-Specially, whether working with languages from the same language family (e.g., English + French) or different language families (e.g., English + Chinese), the results on our dataset were inferior to those obtained by training on a single language.
+Specifically, R1: MLLMs exhibit significant limitations in generalizing to complex dialogue scenarios; R2: Fine-tuning on XMP improves only marginally, with the 70B model achieving at most a 1% absolute gain over its 8B counterpart; R3: The multilingual complementarity effect largely diminishes in complex, multi-party dialogues.
 Our key contributions are summarized as follows:
   • We construct and release XMP, the first high-quality parallel multilingual dataset for multi-party dialogue, covering six languages and over 312k conversations.
-The dataset offers diverse topics and supports research into complex multilingual dialogue generation and cross-lingual transfer.
-  • We conduct extensive evaluations on XMP, revealing significant limitations in the multilingual capabilities of current LLMs, particularly their inability to leverage multilingual complementary effects in complex dialogue settings.
-  • We provide detailed analysis and insights into the mechanisms of LLMs in multilingual environments, highlighting key challenges and opportunities for improving their adaptability to real-world, multi-party dialogue scenarios.
+The dataset offers diverse topics and supports research into multilingual multi-party conversation and cross-lingual transfer.
+  • We conduct extensive evaluations on XMP, revealing significant limitations in the multilingual capabilities of current LLMs.
+  • We reveal key insights into the performance of MLLMs, providing valuable guidance for advancing multilingual dialogue systems in realistic and challenging conversational contexts.
 2 Related Work
 2.1 Multilingual Datasets for Dialogue
-The development of high-quality multilingual dialogue datasets has significantly contributed to the advancement of multilingual research.
-In recent years, multilingual datasets have expanded to cover a wide range of tasks, such as XCOPA Ponti et al.
-(2020), which focuses on multiple-choice question answering, and XWinograd Muennighoff et al.
-(2022), which emphasizes reasoning tasks.
-However, dialogue-related datasets remain scarce and are often constrained to simple dialogue scenarios.
-For instance, XDailyDialog Liu et al.
-(2023) provides a multilingual parallel dialogue dataset with four languages, designed for response retrieval tasks.
-Similarly, XPersona Lin et al.
-(2021) extends the Persona-Chat dataset to six languages, focusing on cross-lingual personalized dialogue.
-While these datasets support tasks such as monolingual and cross-lingual dialogue, they are limited in complexity, often focusing on straightforward conversation patterns and neglecting multi-party and multi-turn dialogue structures.
-Moreover, they tend to overlook the crucial task of dialogue response generation, which is fundamental for building robust dialogue systems.
+High-quality multilingual datasets play a crucial role in advancing multilingual NLP research, such as XCOPA Ponti et al.
+(2020) and XWinograd Muennighoff et al.
+(2022) primarily focus on structured tasks like multiple-choice question answering and reasoning.
+Although there exist multilingual dialogue datasets such as XDailyDialog Liu et al.
+(2023) and XPersona Lin et al.
+(2021), these datasets primarily emphasize straightforward two-party dialogues, typically lack complexity, rarely addressing multi-party scenarios, and frequently neglect the critical task of response generation.
+Given the fundamental importance of generating coherent and contextually relevant dialogue responses, this omission significantly restricts progress in multilingual dialogue system development.
+2.2 Dialogue in Complex Scenarios
+While dialogue generation has been extensively studied, most existing benchmarks focus on dyadic (two-party) conversations, which significantly limits their applicability to more realistic, multi-party scenarios.
+Friends Shmueli and Ku (2019) is one of the most commonly used multi-party dialogue datasets, but it suffers from a key limitation: it includes only six fixed speakers from scripted television show dialogues.
+This restricts its generalizability, as it lacks the diversity and spontaneity found in real-world multi-party interactions.
+Ubuntu Dialogue Corpus (UDC) Lowe et al.
+(2015) is another widely used dataset featuring technical support conversations from the Ubuntu community.
+However, the dialogues in UDC are primarily noisy forum threads with fragmented turn-taking and frequent context gaps.
+Moreover, the conversations are largely constrained to a narrow domain of technical troubleshooting, offering limited topic diversity.
 2.3 Multilingual Capabilities of Large Language Models
-Recent advancements in large language models (LLMs), such as GPT-4 Achiam et al.
-(2023) and LLaMA Touvron et al.
-(2023), have showcased impressive multilingual capabilities.
-By leveraging instruction tuning, human feedback alignment Lai et al.
-(2024), and vast multilingual training corpora, these models have significantly improved performance across diverse languages, including many low-resource ones.
-However, their evaluation has predominantly focused on relatively structured tasks, rather than the more dynamic and context-rich settings of dialogue generation.
-Multilingual complementarity, a key feature of cross-lingual transfer, has been extensively studied in the context of multilingual language models (MLMs).
+Recent multilingual large language models (MLLMs), such as GPT-4 Achiam et al. (2023) and LLaMA Touvron et al. (2023), have shown strong multilingual capabilities Lai et al. (2024).
+Further more, Cross-lingual data sharing has been extensively studied in the context of multilingual language models (MLMs).
 Choenni et al.
-(2023) investigated cross-lingual data sharing during fine-tuning and demonstrated that training samples from one language can reinforce or complement the knowledge acquired from other languages.
+(2023) investigated how training samples from one language can influence the learning of other languages during fine-tuning.
 Similarly, Doddapaneni et al.
 (2021) highlighted that multilingual models often achieve better performance by leveraging linguistic similarities across languages, enabling effective zero-shot and few-shot cross-lingual transfer.
-These findings suggest that MLMs can effectively exploit shared representations across languages for structured tasks, such as natural language inference and multiple-choice question answering.
-However, despite these promising results, the adaptability of large language models (LLMs) to complex dialogue scenarios remains underexplored.
-Most existing research focuses on relatively structured tasks and overlooks dynamic settings like multi-turn, multi-party dialogue.
-Our findings indicate that multilingual complementarity often diminishes in these complex dialogue scenarios.
-For instance, training a model on both English and another language (e.g., French or Chinese) does not consistently outperform single-language training, even for linguistically similar languages.
+Building upon these insights, we introduce the term Multilingual Complementarity to describe the phenomenon where multilingual models leverage data from multiple languages to enhance performance on a specific language task.
+While prior studies have observed this effect in structured tasks, its presence in complex generation tasks remains underexplored.
 3 Preliminary Experiments
-To investigate the capabilities of large language models (LLMs) in generating responses for complex dialogue scenarios, we conducted preliminary experiments using approximately 1,000 English multi-party dialogue samples, each involving three or more speakers discussing topics such as culture, society, and entertainment.
+To investigate the capabilities of large language models (LLMs) in generating responses for complex dialogue scenarios, we conducted preliminary experiments using approximately 1,000 English multi-party dialogue samples.
+The detailed experimental setup and results are provided in Appendix A.
 By testing various templates and different numbers of in-context learning (ICL) examples, we observed that LLMs failed to achieve significant improvements from ICL.
-Even for English, the language LLMs are typically most proficient in, the results showed almost no variation in performance as the number of ICL examples increased, with automatic metrics such as BLEU and ROUGE remaining nearly unchanged.
+Even for English, the language LLMs are typically most proficient in, the results showed almost no variation in performance as the number of ICL examples increased.
 The phenomenon raised doubts about the ability of LLMs to handle complex dialogue scenarios.
@@
 The results revealed a consistent performance decline for all non-English languages as the number of ICL examples increased, regardless of template design.
-The detailed experimental setup and results are provided in Appendix A.
 These findings prompt deeper questions about the nature of LLMs’ capabilities and why ICL fails to yield improvements in this scenario.
@@
 However, in complex multi-party dialogue settings, where multiple speakers interact dynamically and linguistic nuances vary significantly across languages, the models’ reliance on data-driven pattern matching seems insufficient.
-They struggle to incorporate additional examples to refine their understanding of the dialogue’s structure or adapt their reasoning to the specific task context.
+They struggle to incorporate additional examples to refine their understanding of the dialogue’s structure or adapt their reasoning to the specific task context (R1).
 4 Data Collection
-The Multilingual Complex Dialogue Dataset aims to provide high-quality parallel data for studying multilingual in complex multi-party scenarios.
-The raw data was collected from podcasts and subsequently translated into parallel corpora using Google Translate.
-In this section, we describe the process of collecting and preparing the dataset.
+To further evaluate the capabilities of MLLMs in multi-party dialogue and to support future related research, we propose XMP dataset.
 4.1 Data Source
-In our data collection process, we chose podcasts as the primary data source due to their ability to provide high-quality transcripts and their frequent use of multi-party interaction formats involving three or more speakers.
-The format aligns well with the needs of multi-party dialogue research and offers a realistic representation of complex dialogue scenarios found in real-world interactions.
 We collected data from various publicly available podcasts, including "Life Kit," "Pop Culture Happy Hour," "Short Wave," "Throughline," "Consider This from NPR," "It’s Been a Minute," "The NPR Politics Podcast," "The Lawfare Podcast," and "Planet Money." All these programs are publicly accessible and available for free download.
 Details regarding the crawling, filtering, cleaning, and segmentation processes are provided in Appendix B.
+We chose podcasts as the primary data source due to their ability to provide high-quality transcripts and their frequent use of multi-party interaction formats involving three or more speakers.
 4.2 Translation Method
-After completing the initial collection of English transcripts, the next challenge was expanding them into multiple languages without compromising contextual integrity or the complexity of the dialogue scenarios.
-The quality of large-scale parallel corpora depends heavily on the accuracy and coherence of the translation at the word, sentence, and conversation levels.
-To address this, we designed a comprehensive Translation Method Evaluation process to strike an optimal balance between translation quality and efficiency.
-Translation Method W-Level U-Level C-Level
-Google Translate 92.6 86.9 84.5
-Baidu Translate 88.2 73.1 74.7
-DeepL Translate 88.7 70.8 73.4
-iFLYTEK Translate 86.2 70.2 72.3
-LLaMA 3.1 90.4 80.3 82.8
-ChatGPT-4o 94.5 85.9 87.8
-Human Translation 99.2 98.5 95.4
-Table 2: Evaluation Results of Translation Methods
+Singh et al.
+(2024) report that traditional machine translation systems often outperform large language models in translation quality, however, they tend to be less sensitive to cultural nuances.
+To validate which translation strategy is more suitable for our task, we evaluated multiple translation methods.
+Specifically, we experimented with commonly used tools in prior research Liu et al.
+(2023); Lin et al.
+(2021), including Google Translate, DeepL, Baidu Translate, and iFLYTEK Translate, alongside LLaMA 3.1 and ChatGPT-4o.
+We adopted both human and automatic evaluations to ensure translation quality.
+For human evaluation, we selected 50 English dialogue samples spanning diverse topics and lengths (7–15 turns), and translated them into five languages (Chinese, Japanese, French, German, Italian).
+Three linguists reviewed each sample for word-level accuracy, fluency, and dialogue-level coherence.
+ChatGPT-4o and Google Translate received the highest overall ratings, though ChatGPT-4o occasionally introduced off-topic phrases (e.g., “Here is the translation…”), making it less suitable for large-scale batch translation.
+Given its consistency,cost, speed, and prior use in similar work, we selected Google Translate.
+To further support this decision, we translated 5,000 samples into Chinese and German using both GPT-4o and Google Translate, and trained LLaMA 3.1 on each dataset.
+The generation outcomes showed minimal differences, with Google Translate slightly outperforming in some metrics.
+These results reinforce its suitability for scalable multilingual dialogue construction.
+Full evaluation details and translation case analyses are provided in Appendix C
+4.3 Dataset Quality Analysis
+As mentioned before, Singh et al.
+(2024) purpose that machine translation systems tend to be less sensitive to cultural nuances.
+To systematically evaluate the quality of the whole dataset, we conducted a thorough assessment focusing on two key aspects: Cross-lingual Consistency Analysis and Linguistic and Cultural Consistency.
+Dataset Task Type Multilingual Complex Context
+UDC RG No Yes
+XCOPA MCQA Yes No
+XWinograd Reasoning Yes -
+XDailyDialog RR Yes No
+XMP(Ours) RG Yes Yes
+Table 1: Common multilingual or multi-party datasets. Task types include RG for Response Generation, RR for Response Retrieval, and MCQA for Multiple-Choice Question Answering.
 Statistic Train Test Total
@@
 Table 2: Statistics of our dataset.
-We began by selecting 50 representative samples that covered a diverse range of topics (e.g., law, politics, daily life), dialogue turns (ranging from 7 to 15 turns), and multi-party podcast segments.
-This ensured that the evaluation samples were sufficiently diverse in complexity, providing reliable references for subsequent large-scale translation.
-For these samples, we employed several translation methods, including Google Translate, DeepL, Baidu Translate, iFLYTEK Translate, LLaMA 3.1, ChatGPT-4o and human translation.
-Each method translated the English transcripts into five other languages (Chinese, Japanese, French, German, and Italian), resulting in 300 translations (50 samples × 6 languages).
-To ensure objectivity and reliability, three linguistics experts reviewed each translation in multiple rounds.
-The review process identified issues such as lexical misusage, semantic distortion, or contextual inconsistency in the translated outputs.
-Ultimately, the evaluation focused on three dimensions:
-  • Word-level evaluation assessed the accuracy of individual word translations and checked for spelling errors.
-  • Utterance-level evaluation examined sentence fluency, semantic accuracy, and appropriateness in context.
-  • Conversation-leve evaluation ensured contextual consistency and alignment with the original dataset.
-As shown in Table 2, human translation consistently achieved the highest overall quality.
-However, due to the prohibitive cost and extended timeline of large-scale manual translation, we explored automated tools.
-Among Google Translate, DeepL, Baidu, iFLYTEK, LLaMA 3, ChatGPt-4o, both ChatGPT-4 and Google Translate have shown impressive results, with ChatGPT-4 slightly outperforming Google Translate.
-However, ChatGPT-4 requires careful handling of templates, as it may provide natural translations for some examples but sometimes offers literal translations that overlook specific contexts.
-Considering translation speed and cost-effectiveness, we opted for Google Translate as the more practical choice for large-scale multilingual tasks.
-Specific case comparisons of the translation tools are detailed in Appendix C.
-4.3 Dataset Quality Analysis and Statistics
-The dataset includes six languages: English, Chinese, Japanese, French, German, and Italian.
-While English serves as the original source, all other languages were generated using translation API.
-To systematically evaluate the quality of the dataset, we conducted a thorough assessment focusing on four key aspects: word-level accuracy, sentence fluency, semantic consistency, and contextual coherence.
-To assess the quality of the dataset, we employed expert evaluation.
-Specially, we randomly selected 500 conversations, covering various topics.
+Specially, we randomly selected 1000 conversations (200 samples for each language), covering various topics.
 For each language, two evaluation experts were selected based on their proficiency in both English and the target language.
@@
 Detailed evaluation criteria and the scoring results for each language are provided in Appendix D.
+Overall, we found that the dataset demonstrated good quality across the board, achieving strong results in both Cross-lingual Consistency and Linguistic and Cultural Consistency, indicating that the translated texts can preserve semantic consistency while also taking cultural differences into account to some extent.
+Although relatively low-resource languages such as Italian received slightly lower scores than others, their performance still fell within an acceptable range.
+4.4 Dataset Statistics
+In Table 1, we compare our XMP dataset with several commonly used multilingual or multi-party datasets. XMP is the first high-quality parallel multilingual dataset for multi-party dialogue.
 Figure 2: Distribution of Speaker Counts
-In Table 3, we present the detailed statistical analysis of the dataset.
-The dataset contains a total of 312k dialogues and 2,649k utterances, with an average of 8.47 turns per dialogue, indicating a high level of dialogue density.
-This ensures that each dialogue typically involves multiple exchanges, creating scenarios rich in multi-turn interactions.
-The dataset also emphasizes its multi-party characteristics, with an average of 3.64 speakers per dialogue.
-Notably, over 75% of the dialogues involve at least three speakers, as shown in Figure 1, reflecting the interaction between multiple distinct roles.
-This multi-party nature enhances the realism of the dataset and provides a robust foundation for studying complex dialogue scenarios.
-Moreover, the average number of turns per speaker is 2.33, indicating that most speakers contribute multiple times to the conversation.
+In Table 2, we present key statistics of the dataset, which includes 312k dialogues and 2,649k utterances, averaging 8.47 turns per dialogue.
+This high dialogue density supports rich multi-turn interactions.
+On average, each dialogue features 3.64 speakers, and over 75% involve at least three participants (Figure 2), highlighting the dataset’s multi-party nature and the diverse role dynamics it captures.
+Each speaker contributes an average of 2.33 turns, suggesting active participation across the dialogue.
 Additionally, we ensured that the target speaker (i.e., the speaker for whom a response is to be generated) had spoken at least once earlier in the dialogue history.
@@
 These models were chosen for their demonstrated capabilities in multilingual dialogue response task.
-Their inclusion allows us to examine the trade-offs between model size, and response quality across six languages
+Their inclusion allows us to examine the trade-offs between model size, and response quality across six languages.
 5.2 Task Setup
 The generation task focuses on producing high-quality responses in complex multi-party dialogue scenarios.
-Given a conversational context and the current speaker, the model is tasked with generating a coherent, contextually relevant, and linguistically accurate response.
-Specially, Let $D={(H,S,R)}^{n}_{i=1}$represent the multilingual dialogue dataset for complex scenarios, where $H$ is the dialogue history composed of multiple “Speaker: Utterance” pairs concatenated sequentially, $S$ is the current speaker, and $R$ is the ground truth response.
+Specially, Let $D={(H,S,R)}^{n}_{i=1}$ represents the multilingual dialogue dataset for complex scenarios, where $H$ is the dialogue history composed of multiple “Speaker: Utterance” pairs concatenated sequentially, $S$ is the current speaker, and $R$ is the ground truth response.
 To guide the model in generating speaker-specific responses, a prompt is dynamically constructed based on the current speaker, such as “What will Speaker-x say next?”.
@@
 5.3 Implementation Details
-We utilized LLaMAFactory as the training framework and applied LoRA (Low-Rank Adaptation) for efficient fine-tuning.
+We utilized LLaMAFactory as the training framework and applied LoRA (Low-Rank Adaptation) Hu et al.
+(2021) for efficient fine-tuning.
 The training was conducted on NVIDIA A100 GPUs using the AdamW optimizer with a learning rate of $1e^{-5}$, a warm-up of 500 steps, and a batch size of 8.
@@
 The results highlight the over-reliance of LLMs on high-resource languages.
-The performance gain of LLaMA-3.1 70B over smaller models (e.g., 8B or Qwen-2.5 7B) is marginal, with improvements typically within 1% across most metrics.
+The performance gain of LLaMA-3.1 70B over smaller models (e.g., 8B or Qwen-2.5 7B) is marginal, with improvements typically within 1% across most metrics (R2).
 This suggests that the emergent abilities often associated with larger models are not prominently reflected in this task.
 Instead, the slight advantage may stem from the larger models’ capacity to generate grammatically and lexically refined text, rather than demonstrating a better understanding of the task.
-Additionally, Japanese yields unexpectedly high scores across all models.
-Despite its low proportion in pretraining corpora, Japanese outperforms other non-English languages.
-We hypothesize that this stems from the unique syntactic structure and word composition of Japanese, which relies heavily on fixed sentence patterns and the use of kana.
-Tokenization using Janome likely also contributes to higher n-gram overlap, boosting surface-level metrics like BLEU and ROUGE.
-To validate this hypothesis, we included BERTScore as an additional evaluation metric.
-Unlike n-gram metrics, BERTScore evaluates semantic similarity.
-6.2 Semantic Analysis with BERTScore
+6.2 Structural and Semantic divergence
+Japanese and Chinese both achieve relatively high scores on surface-level metrics like BLEU and ROUGE, Japanese even outperforming all other languages besides English, and Chinese ranking among the top as well.
+This can be attributed to factors such as syntactic regularity, frequent use of fixed expressions and kana in Japanese, and strong n-gram overlaps that benefit token-based evaluation.
+However, as shown in Table 4, BERTScore tells a different story: Japanese aligns semantically with other non-English languages, while Chinese records the lowest score (0.6086), indicating significant structural and semantic divergence that models struggle to capture.
 Language BERTScore
@@
 Table 4: BERTScore results for LLaMA-3.1 8B across different languages.
-Table 4 presents BERTScore results for all six languages.
-English achieves the highest semantic alignment, confirming its strong representation in training data.
-Japanese, despite its high BLEU and ROUGE scores, shows comparable BERTScore results to other non-English languages, indicating that its apparent advantage in traditional metrics stems from tokenization effects rather than true semantic superiority.
-For other languages, the results indicate that Japanese achieves a comparable BERTScore to other non-English languages, suggesting that its seemingly superior performance in traditional metrics reflects tokenization effects rather than genuine semantic alignment.
-In contrast, Chinese scores similarly to other non-English languages in metrics such as F1, BLEU, ROUGE-1, ROUGE-2, and ROUGE-L, yet records the lowest BERTScore at just 0.6086.
-This discrepancy indicates significant differences in structural and semantic expression between Chinese and other languages.
-Unlike many other languages, Chinese has less rigid grammatical requirements, allowing sentences with entirely different meanings to use similar words and word orders.
 6.3 Multilingual Complementary Ability
@@
 6.3.1 Mixed Training with English
-Previous research often highlights that mixing English as a source and target language during training provides significant benefits for the target language due to English’s dominance in pretraining corpora and its rich semantic structure.
+Previous researches Whitehouse et al.
+(2023); Li et al.
+(2024) often highlight that mixing English as a source and target language during training provides significant benefits for the target language due to English’s dominance in pretraining corpora and its rich semantic structure.
 However, our experiments on XMP challenge this assumption.
@@
 This suggests that while larger model capacities can temper the negative impacts of mixed training, the benefits derived from cross-lingual interaction are marginal and insufficient to fully leverage the potential of multilingual complementary ability in complex dialogue scenarios.
+(R3)
 6.3.2 Cross-Language Benefits Are Asymmetrical
 The experimental results for 70B-model reveal an intriguing asymmetry: while English typically fails to enhance the performance of other languages, other languages, such as French or German, often provide improvements when mixed with English.
-It might stem from the model’s reliance on English during pretraining, making it more adept at leveraging additional multilingual context for English-centric tasks while struggling to adapt English’s resources to enhance other languages effectively.
 6.3.3 Effect of Linguistic Similarity
-The relationship between linguistic similarity and cross-lingual gains appears to be more complex than previously assumed.
-Traditionally, languages within the same linguistic family, such as English and German, or French and English, are expected to exhibit significant cross-lingual transfer benefits when trained together.
+Both Eronen et al.
+(2023) and Aepli and Sennrich (2022) report that linguistic similarity plays a crucial role in cross-lingual transfer.
+Eronen et al.
+(2023) demonstrate this across sentiment analysis, and dependency parsing.
+Similarly, Aepli and Sennrich (2022) observe that in POS tagging, transfer between closely related language varieties yields significantly better performance than between unrelated pairs.
 However, our experimental findings suggest that these linguistic connections do not guarantee consistent or significant cross-lingual improvements.
@@
 The nuanced relationship highlights the need for deeper investigations into how linguistic similarity interacts with other variables in multilingual training, particularly in complex and dynamic tasks.
-It underscores that shared linguistic heritage, while beneficial, is not a deterministic factor for cross-lingual success, and the benefits of multilingual training are far from uniform across different language pairs
+It underscores that shared linguistic heritage, while beneficial, is not a deterministic factor for cross-lingual success, and the benefits of multilingual training are far from uniform across different language pairs.
+6.4 Probing Context Sensitivity
+Figure 3: Case Study
+To examine how effectively the model leverages extended dialogue history, we conducted a controlled evaluation by randomly removing 30% of the context while retaining the final two turns.
+The rationale behind this setup is that if the model truly integrates long-range contextual information, removing part of the earlier context should noticeably impact its performance.
+We applied this evaluation to the LLaMA 3 8B model on both English and Chinese subsets. Results are presented in Table 6. Overall, the differences across both languages were minimal.
+These findings suggest that the model does not effectively utilize broader dialogue history.
+Its performance appears to rely primarily on recent turns, reflecting a limited capacity to model longer-range conversational dependencies.
+Language Context F1 B-1 B-2 R-L
+Chinese Full 18.37 11.60 4.68 5.44
+Chinese - 30% 18.52 12.19 4.77 5.85
+English Full 19.92 12.80 5.73 5.54
+English - 30% 19.36 12.60 5.41 4.91
+Table 6: Performance comparison with and without 30% context removal on the LLaMA 3 8B model. Context removal excludes the final two turns.
+6.5 Case Study
+To more intuitively demonstrate the current limitations of large language models in understanding multi-party conversations, we conduct a case study and analyze the model’s predictions.
+We found that the model’s responses often: 1.
+Misinterpreted conversational context and overlooked the roles of specific participants.
+2.
+Lacked concrete factual content
+We presented several cases in the Figure 3 . Due to character limitations, all examples shown here are in English — the model’s performance tends to degrade further in other languages.
+In the first case, the model ignored the speaker Harris and abruptly shifted to a discussion of Biden’s plan.
+In contrast, the ground truth continued with a discussion of Harris’ media strategy, which was contextually appropriate.
+We attribute this error to amplified hallucination under complex dialogue settings.
+In the second example, the model predicted a response like “I hope so,” which contributed nothing meaningful to the conversation, whereas the ground truth was a natural continuation of the ongoing topic.
+We believe this failure is likely due to: 1.
+the model struggles to maintain awareness of the broader conversational flow and context, particularly in multi-party settings; 2.
+current large models’ pretraining corpora still lack rich, multi-turn, real-world conversational data.
 7 Conclusion
-In this study, we introduced XMP, a high-quality multilingual parallel multi-party dialogue dataset, designed to address the gaps in existing datasets for complex dialogue scenarios.
-XMP provides extensive coverage of six languages to address the lack of comprehensive resources for complex multilingual dialogue generation.
-Using XMP as a benchmark, we evaluated the multilingual complementary ability of LLaMA-3.1 8B and 70B models, uncovering unexpected findings that challenge prior research.
-Mixing English with other languages during training often resulted in performance degradation, particularly for smaller models, while larger models like the 70B variant demonstrated better resilience and achieved marginal gains in specific cases, such as English-French mixing.
-Interestingly, linguistic similarity did not always yield mutual reinforcement; for instance, English and German showed negative impacts when mixed, whereas French consistently enhanced performance when paired with other languages.
-These insights underscore the limitations of current multilingual models and highlight the need for further research into linguistic properties and cross-lingual transfer, with XMP serving as a robust benchmark for advancing multilingual dialogue generation.
+This study aims to explore the generalization capabilities of MLLM in complex dialogue scenarios.
+To enable this investigation, we construct XMP, a high-quality, parallel multilingual multi-party dialogue dataset.
+XMP serves as a critical resource to support rigorous evaluation in this challenging setting.
+Through extensive experiments, we find: R1: MLLMs exhibit significant limitations in generalizing to multi-party dialogue.
+R2: Fine-tuning on XMP yields only marginal improvements, with larger models showing limited gains.
+R3: Contrary to prior studies, multilingual complementarity effects are inconsistent or even detrimental.
+We hope our work and the XMP dataset will inspire further efforts toward building more context-aware, robust, and multilingual conversational AI.
 8 Limitations
@@
 To investigate this further, we expanded the 1,000 English samples to include French, German, Italian, and Chinese.
+The results are shown in Figure 5.
 Figure 4: This is an case with the target language in Chinese, showcasing three different template types (Type 1, Type 2, Type 3).
 Figure 5: The results of Preliminary Experiments. $\{metric\}\_{n}$ means n examples used in the ICL and $\{Language\}\_{x}$ means Type X template used.
-Following the approach of CCD et al., we applied different templates combined with ICL for further validation.
+Following the approach of Choenni et al.
+(2023) et al., we applied different templates combined with ICL for further validation.
 Specially, we designed three different experimental methods to validate the model’s performance variations:
@@
 iFlytek {CJK}UTF8gbsn一位未经验考验的制片人和他的一帮完全不知名的演员准备创造电视历史,而电视台的西装束紧着他们。几乎所有可能出错的事情会发生,就像,完全,彻底。这很糟糕。我是艾莎·哈里斯。
-LLaMA3 {CJK}UTF8gbsn一位毫无经验的制片人和他全新演员阵容准备创造电视史,可电视台的高层却紧盯着他们。几乎所有可能出错的地方都出了问题,真是糟透了。我是艾莎·哈里斯。
-Human {CJK}UTF8gbsn一位未尝试的制片人带领一群毫无名气的演员,准备改写电视规则,可电视台高层却紧盯着他们的一举一动。结果几乎出错的地方都有错,真是糟糕了。我是艾莎·哈里斯。
+LLaMA3 {CJK}UTF8gbsn一位毫无经验的制片人和他全新演员阵容准备创造电视史,可电视台的高层却紧盯着他们。几乎所有可能出错的地方都会发生,真是糟透了。我是艾莎·哈里斯。
+Human {CJK}UTF8gbsn一个毫无经验的制片人带领一群毫无名气的演员,准备改写电影史,可电视台高层却死死盯着他们的一举一动。结果几乎所有地方都出差错了,真是糟透了。我是艾莎·哈里斯。
 Table 7: Specific case comparisons of the translation tools.
@@
     Response Quality: Segments with overly short or long responses, or responses deemed meaningless (e.g., “Yeah, wel” or “Alright”), were excluded.
-Appendix C Chosen of Translate API
+Appendix C Comparison of Translate Method
+C.1 Human Evaluation for Different Translation Methods
+Translation Method W-Level U-Level C-Level
+Google Translate 92.6 86.9 84.5
+Baidu Translate 88.2 73.1 74.7
+DeepL Translate 88.7 70.8 73.4
+iFLYTEK Translate 86.2 70.2 72.3
+LLaMA 3.1 90.4 80.3 82.8
+ChatGPT-4o 94.5 85.9 87.8
+Human Translation 99.2 98.5 95.4
+Table 8: Evaluation Results of Translation Methods. W-Level, U-Level and C-Level refer to Word-Level, Utterance-Level and Conversation-Level.
+Language Method F1 B-1 B-2 R-L
+Chinese GPT-4o 17.45 10.93 4.59 15.01
+Chinese Google 17.61 11.01 4.55 14.97
+German GPT-4o 16.97 11.60 4.91 13.88
+German Google 17.12 11.51 4.95 13.84
+Table 9: Performance comparison of LLaMA 3.1 trained on translations generated by GPT-4o vs. Google Translate (5k samples each).
+For human evaluation, we began by selecting 50 representative samples that covered a diverse range of topics (e.g., law, politics, daily life), dialogue turns (ranging from 7 to 15 turns), and multi-party podcast segments.
+Each method translated the English transcripts into five other languages (Chinese, Japanese, French, German, and Italian), resulting in 300 translations.
+To ensure objectivity and reliability, three linguistics experts reviewed each translation in multiple rounds.
+The review process identified issues such as lexical misusage, semantic distortion, or contextual inconsistency in the translated outputs.
+Ultimately, the evaluation focused on three dimensions:
+  • Word-level evaluation assessed the accuracy of individual word translations and checked for spelling errors.
+  • Utterance-level evaluation examined sentence fluency, semantic accuracy, and appropriateness in context.
+  • Conversation-level evaluation ensured contextual consistency and alignment with the original dataset.
+As shown in Table 8, among Google Translate, DeepL, Baidu, iFLYTEK, LLaMA 3, ChatGPt-4o, both ChatGPT-4 and Google Translate have shown impressive results, with ChatGPT-4o slightly outperforming Google Translate.
+However, ChatGPT-4o requires careful handling of templates, as it may provide natural translations for some examples but sometimes offers literal translations that overlook specific contexts.
+Considering translation speed and cost-effectiveness, we opted for Google Translate as the more practical choice for large-scale multilingual tasks, just like prior works Lai et al.
+(2024); Liu et al.
+(2023).
+C.2 Automatic Evaluation for Different Translation Methods
+We further validated translation quality through a small-scale generation experiment, as shown in Table 9.
+A subset of approximately 5,000 English samples was translated into Chinese and German using both GPT-4o and Google Translate.
+We then fine-tuned LLaMA 3.1 using each translated dataset and compared dialogue generation performance.
+Results showed minimal performance differences between the two translation sources.
+These findings support the use of Google Translate as a practical and reliable solution for large-scale multilingual dialogue construction.
+C.3 Case comparisons
+Specific case comparisons of the translation tools are detailed below.
 Origin Prompt 1:{CJK}UTF8gbsn翻译成中文 Prompt 2:{CJK}UTF8gbsn翻译成中文,请不要直译,在这句话中有很多需要理解并意译的地方,请注意
@@
 The result with prompt 2 is much better than result with prompt 1.
-To further illustrate our reasoning for choosing the Google Translate API, the Table 7 and Table C present an actual translation snippet of a dialogue.
+To further illustrate our reasoning for choosing the Google Translate API, the Table 7 and Table C.3 present an actual translation snippet of a dialogue.
 In terms of word-level evaluation, all the translation APIs demonstrated accurate spelling without errors.
@@
 To assess the quality of the dataset, we employed expert evaluation.
-Specifically, we randomly selected 500 conversations, covering various topics.
+Specifically, we randomly selected 1000 conversations (200 samples for each language), covering various topics.
 The review was conducted based on the following aspects, with scores ranging from 1 to 5 for each metric.
-The final scores and detailed scoring criteria are shown in Table 9.
-Metric Score Range Description Final Score
-Conversation Diversity 1 Very Low Diversity: Narrow range of topics, repetitive dialogue structures, limited interaction patterns. 4.74
-2 Low Diversity: Few different topics, minimal variation in interaction styles, some repetition.
-3 Moderate Diversity: A reasonable range of topics, some diversity in interaction styles and contexts.
-4 High Diversity: Diverse topics and contexts, multiple types of dialogues and varied interactions.
-5 Very High Diversity: Rich variety of topics, multiple communication styles, diverse contexts and subjects.
-Conversation Fluency 1 Very Low Fluency: Disjointed conversations with abrupt topic changes, frequent grammatical errors. 4.21
-2 Low Fluency: Awkward transitions, slight disruptions in flow, minor grammatical issues.
-3 Moderate Fluency: Clear dialogue with occasional awkward phrasing, understandable overall.
-4 High Fluency: Natural flow, clear transitions between topics, minor errors not affecting overall fluency.
-5 Very High Fluency: Seamless conversation with smooth transitions, grammatically correct throughout.
-Annotation Quality 1 Very Low Quality: Responses are either too short or excessively long, often irrelevant or nonsensical.
-Responses frequently deviate from the previous context, making them incoherent or disjointed.
-4.43
-2 Low Quality: Responses are occasionally too short or too long, with some parts being irrelevant.
-There are noticeable inconsistencies with the previous context, leading to slight confusion in the conversation flow.
-3 Moderate Quality: Responses are generally of appropriate length but may occasionally lack relevance or slightly stray from the context.
-They remain understandable and contextually coherent, with some minor lapses.
-4 High Quality: Responses are of reasonable length, relevant, and mostly meaningful.
-There is good consistency with the previous dialogue, but small deviations or slight misunderstandings may still occur.
-5 Excellent Quality: Responses are always appropriately lengthened, coherent, and directly related to the previous dialogue.
-They are contextually meaningful, contributing effectively to the conversation without unnecessary repetition or confusion.
-Table 9: Detailed scoring criteria for dataset quality evaluation.
+Cross-lingual Consistency Analysis
+use the Scoring Criteria below to assess how well the target-language text preserves the semantic meaning and conceptual structure of the English source text.
+  • Semantic Equivalence: Accuracy in capturing key information, terminology, and meaning from the source.
+  • Conceptual Alignment: Faithfulness to the overall theme, reasoning chain, and logical flow of the source text.
+The Scoring Criteria and results are shown in Table 11 and Table 12.
+Linguistic and Cultural Consistency
+focuses on the quality of the target-language text in its own linguistic and cultural context.
+  • Naturalness: Whether sentences are fluent, grammatically sound, and idiomatic.
+  • Cultural Appropriateness: Alignment with cultural norms, conventions, and stylistic preferences.
+The Scoring Criteria and result are shown in Table 13 and Table 14.
+Score Semantic Equivalence Conceptual Alignment
+1 Many key details lost or mistranslated; Significant deviation from source meaning Main topic is unclear or completely misaligned; Severe logical gaps or missing core concepts
+2 Substantial inaccuracies in important information; Inadequate handling of key terms Overall theme is partially discernible, but reasoning steps are shaky; Noticeable gaps in coherence
+3 Primary content is mostly correct; Some inaccuracies in less crucial details Main idea is largely maintained; Minor logical jumps but generally comprehensible
+4 Nearly all key information is accurately preserved; Only minor misinterpretations The text closely follows the source theme; Reasoning and structure are coherent, with few gaps
+5 All core information perfectly conveyed; No noticeable semantic errors Fully aligned with source logic; Clear and consistent reasoning; Nearly flawless
+Table 11: Semantic Equivalence and Conceptual Alignment Scoring Guidelines
+Target Language Samples Semantic Equivalence Conceptual Alignment
+ZH (Chinese) 200 4.13 4.02
+JA (Japanese) 200 4.17 4.11
+FR (French) 200 4.32 4.25
+DE (German) 200 4.05 3.92
+IT (Italian) 200 3.93 3.88
+Table 12: Quality Scores by Language
+Score Naturalness Cultural Appropriateness
+1 Numerous grammatical or lexical errors. Difficult to read or parse Severely mismatched with target-language cultural norms. Potentially offensive or highly unnatural expressions
+2 Limited fluency; awkward or clumsy phrasing; Noticeable "translationese" Certain mismatches in cultural context; Expressions that appear slightly off or inauthentic
+3 Mostly coherent sentences; Minor issues with word choice or syntax Understandable but not fully adapted to local norms; Some cultural references may be vague or misplaced
+4 Generally fluent and easy to follow; Few or negligible foreign constructs Largely compatible with the target culture; Stylistically appropriate in most contexts
+5 Exceptionally natural, no apparent translation artifacts; Flawless grammar Perfect cultural fit; expressions and style feel native; No inaccuracies or incongruities in cultural references
+Table 13: Scoring Criteria for Naturalness and Cultural Appropriateness
+Target Language #Samples Naturalness Cultural Appropriateness
+ZH (Chinese) 200 3.94 4.10
+JA (Japanese) 200 4.03 4.12
+FR (French) 200 4.25 4.37
+DE (German) 200 3.88 3.96
+IT (Italian) 200 3.72 3.85
+Table 14: Naturalness and Cultural Appropriateness Scores by Language
 Appendix E Detailed training and inference command